40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Thu Apr 18 19:54:01 2013, tara, Notes, DAQ, EPICS channel for slow feedback to laser C3PSL_LASER_SLOW.png
    Reply  Mon Apr 22 17:13:24 2013, tara, Notes, DAQ, slow feedback to laser via EPICS is on KP.jpgKI.jpgCTN_DAQ_wiring.png
       Reply  Wed May 8 01:11:09 2013, tara, Notes, DAQ, slow feedback to laser via EPICS is on 
Message ID: 1155     Entry time: Mon Apr 22 17:13:24 2013     In reply to: 1151     Reply to this: 1172
Author: tara 
Type: Notes 
Category: DAQ 
Subject: slow feedback to laser via EPICS is on 

The medm screen for the 1st laser is completed, the servo is on an stable. Refcav has been locked for a few hours as of now.

  • The output for slow feedback is on J9, slot 9/10 (VMIVME 4116 C2 S3). This is an unused channel previously assigned to PSL-ISS_ISET, I checked ISS.db file to look for the VMIVME address. For the slot number, I look up the channel name in D980535-C-C document.
  • I added low pass filters (~100mHz) to both input and outputs of EPICS.
  • All EPICS channels for slow feedback and perl scripts are in SLOW_LASER.db file.
  • The startup.cmd file is updated accordingly.
  • servo gain is optimized. ??? What does that mean??? How about some performance plots? (About the bode plot, I'm trying to get a transfer function of the NPRO slow input, with that I can estimate the bode plot of the loop. As of now I just adjust the PID gain so that the loop is stable)

KP.jpg

fig1: FAST feedback to the laser is shown in blue plot, vertical axis:1V/div, horizontal axis:  4 sec/ div. I adjusted proportional gain first, to get only a few overshoots with acceptable rise time.

KI.jpg

fig2: Then I adjusted integral gain to eliminate the offset, and Derivative gain to reduce overshoot. More about PID gain can be found here. Current Value KP = -0.0002, KI =-0.00015, KD = 0.

I set the output to be between -2 V to 9 V. Since we need to lock it to GYRO later, it has to be able to be tuned to match the gyro laser. Currently, Gyro laser is operated around 35 Deg C which is similar to 8V input to slow feedback.

 

I'm trying to draw a cartoon for DAQ wiring in CTN lab for future reference. This is what I have so far. I'll add it in WIKI page.

CTN_DAQ_wiring.png

ELOG V3.1.3-